2. Sodium salts of phosphoric acid, H3PO4, are used extensively in the preparation of buffers. The Ka values of phosphoric acid are 7.5 x 10-3, 6.2 x 10-8, and 4.8 x 10-13. Calculate the corresponding pKa values (3). If you want the buffer to have a pH of 7.4, which of the salts do you need and why ? (2) Note : you do not need to calculate any amounts!
Answer #1

pKa = -logKa

H3PO4/NaH2PO4   Ka1 =7.5x10-3 ; pKa = 2.1

NaH2PO4/Na2HPO4 Ka2 =6.2x10-8 ; pKa = 7.2

Na2HPO4/Na3PO4 Ka3 =4.8x10-13 ; pKa = 12.3

Buffers works best when the pKa corresponds to the desired pH.

Look up the pKa (or Ka) values associated with each acid/base conjugate pair which is closed to the desired pH = 7.4

NaH2PO4/Na2HPO4 is the right choice for the buffer of desired pH.